# Theoretical Insights into Overparameterized Models in Multi-Task and Replay-Based Continual Learning

### v1

@@ C.2 Proof of Theorems
underparameterized regime. In the underparameterized regime where $n_{t}\geq p+2$, the signle-task learner is the unique solution to the Equation 4 and is obtained by $w_{t}=(X_{t}X_{t}^{T})^{-1}X_{t}y_{t}$ according to Lemma C.5. Therefore, we must have
$w_{t}=(X_{t}X_{t}^{T})^{-1}X_{t}y_{t}=(X_{t}X_{t}^{T})^{-1}X_{t}(X^{T}_{t}w^{*}_{t}+z_{t})=w^{*}_{t}+(X_{t}X_{t}^{T})^{-1}X_{t}z_{t}$
$\Rightarrow\mathbb{E}[\mathcal{L}_{t}(w)]=\mathbb{E}[\|w_{t}-w^{*}_{t}\|^{2}]=\mathbb{E}[\|(X_{t}X_{t}^{T})^{-1}X_{t}z_{t}\|^{2}]$
@@
$\mathbb{E}[\mathcal{L}_{t}(w)]$ $=\mathbb{E}[z^{T}_{t}X^{T}_{t}(X_{t}X_{t}^{T})^{-1}(X_{t}X_{t}^{T})^{-1}X_{t}z_{t}]$
$=\sigma^{2}\mathbb{E}[\operatorname{tr}(X^{T}_{t}(X_{t}X_{t}^{T})^{-2}X_{t})]$
$=\sigma^{2}\mathbb{E}[\operatorname{tr}((X_{t}X_{t}^{T})^{-2}X_{t}X^{T}_{t})]$
$=\sigma^{2}\mathbb{E}[\operatorname{tr}((X_{t}X_{t}^{T})^{-1})]$
$=\sigma^{2}\operatorname{tr}(\mathbb{E}[(X_{t}X_{t}^{T})^{-1}]))$
@@
where the two last line comes from the fact that $(X_{t}X_{t}^{T})^{-1}$ follows the inverse-Wishart distribution as $\mathcal{IW}(I_{p\times p},n)$.
overparameterized regime. In the overparameterized regime where $p\geq n_{t}+2$, we look for the solution of optimization 5. Therefore, based on Lemma C.6, we can have
@@
$=\mathbb{E}[\|(I-P_{t})w^{*}_{t}\|^{2}]+\mathbb{E}[z^{T}_{t}(X^{T}_{t}X_{t})^{-1}z_{t}]\qquad\quad(\mathbb{E}[z_{t}]=0)$
$=\mathbb{E}[\|(I-P_{t})w^{*}_{t}\|^{2}]+\frac{n_{t}\sigma^{2}}{p-n_{t}-1}\qquad\qquad\qquad((X^{T}_{t}X_{t})^{-1}\sim\mathcal{IW}(I_{n_{t}\times n_{t}},p))$
@@
Based on Lemma C.6 we can write
$w_{1T}=X_{1T}(X^{T}_{1T}X_{1T})^{-1}y_{1T}$
@@
$\mathbb{E}[\mathcal{L}_{i}(w_{1T})]=\mathbb{E}[\|w_{1T}-w^{*}_{i}\|^{2}]$
$=\mathbb{E}[\|\sum_{s=1}^{T}P_{0s}w^{*}_{s}+X_{1T}^{\dagger}z_{1T}-w^{*}_{i}\|^{2}]$
$=\mathbb{E}[\|\sum_{s=1}^{T}P_{0s}(w^{*}_{s}-w_{i}^{*})+(I-P_{1T})w_{i}^{*}+X_{1T}^{\dagger}z_{1T}\|^{2}]$
$=\mathbb{E}[\|\sum_{s=1}^{T}P_{0s}(w^{*}_{s}-w_{i}^{*})+(I-P_{1T})w_{i}^{*}\|^{2}]+\mathbb{E}[\|X_{1T}^{\dagger}z_{1T}\|^{2}]\qquad\quad(z_{1T}\operatorname{\perp\!\!\!\perp}X_{1T},\mathbb{E}[z_{1T}]=0)$
$=\mathbb{E}[\|\sum_{s=1}^{T}P_{0s}(w^{*}_{s}-w_{i}^{*})+(I-P_{1T})w_{i}^{*}\|^{2}]+\sigma^{2}\operatorname{tr}(\mathbb{E}[(X_{1T}X_{1T})^{-1}])\qquad(z_{1T}\sim\mathcal{N}(0,\sigma^{2}I))$
$=\mathbb{E}[\|\sum_{s=1}^{T}P_{0s}(w^{*}_{s}-w_{i}^{*})+(I-P_{1T})w_{i}^{*}\|^{2}]+\frac{\bar{n}\sigma^{2}}{p-\bar{n}-1}\quad\quad((X_{1T}X_{1T})^{-1}\sim\mathcal{IW}(I_{\bar{n}\times\bar{n}},p))$
@@
$\mathbb{E}[\|\sum_{s=1}^{T}P_{0s}(w^{*}_{s}-w_{i}^{*})+(I-P_{1T})w_{i}^{*}\|^{2}]$
$=\mathbb{E}[\|\sum_{s=1}^{T}P_{0s}(w^{*}_{s}-w_{i}^{*})\|^{2}]+\mathbb{E}[\|(I-P_{1T})w_{i}^{*}\|^{2}]$
@@
$=\sum_{s=1}^{T}\mathbb{E}[\|P_{0s}(w^{*}_{s}-w_{i}^{*})\|^{2}]+\sum_{s=1}^{T}\sum_{\begin{subarray}{c}s^{\prime}=1\\ s^{\prime}\neq s\end{subarray}}^{T}\mathbb{E}[(w^{*}_{s}-w_{i}^{*})^{T}P^{T}_{0s}P_{0s^{\prime}}(w^{*}_{s^{\prime}}-w_{i}^{*})]+(1-\frac{\bar{n}}{p})\|w_{i}^{*}\|^{2}$
Since the distribution of $P_{0s}$ is invariant to the permutation of columns of $X_{1T}$, we can focus on finding $\mathbb{E}[\|P_{01}(w^{*}_{1}-w_{i}^{*})\|^{2}]$ and $\mathbb{E}[(w^{*}_{1}-w_{i}^{*})^{T}P^{T}_{01}P_{0T}(w^{*}_{T}-w_{i}^{*})]$ without loss of generality. In Lemmas C.7 and C.8, we have already calculated similar quantities. Therefore,
$\mathbb{E}[\|P_{01}(w^{*}_{1}-w_{i}^{*})\|^{2}]=\frac{n_{1}}{p}(1+\frac{\bar{n}-n_{1}}{p-\bar{n}-1})\|w^{*}_{1}-w_{i}^{*}\|^{2}$
@@
$\mathbb{E}[(w^{*}_{1}-w_{i}^{*})^{T}P^{T}_{01}P_{0T}(w^{*}_{T}-w_{i}^{*})]=-\frac{n_{T}n_{1}}{p(p-\bar{n}-1)}\langle w^{*}_{1}-w_{i}^{*},w^{*}_{T}-w_{i}^{*}\rangle$
@@
$\mathbb{E}[G(w_{1T})]$ $=\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}[\mathcal{L}_{i}(w_{1T})]$
@@
$\mathbb{E}[K(w_{1T})]$ $=\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}[\mathcal{L}_{i}(w_{i})]-\mathbb{E}[\mathcal{L}_{i}(w_{1T})]$
@@
Next, let $\sigma=0$. Assume the equal number of samples $n_{1}=n_{2}=...=n$ and the same task norm for all tasks $\|w_{1}^{*}\|^{2}=...=\|w_{T}^{*}\|^{2}$. Also assume $\langle w^{*}_{s},w_{i}^{*}\rangle=\|w_{1}^{*}\|^{2}\cos\theta$ for all pair of tasks. The knowledge transfer then simplifies to:
$\mathbb{E}[K(w_{1T})]=\frac{2}{T}\frac{n}{p}(1+\frac{\frac{T}{2}n}{p-Tn-1})\sum_{i=1}^{T}\sum_{s=1}^{T}\langle w^{*}_{s},w_{i}^{*}\rangle-\frac{n}{p}(1+\frac{1}{T}+\frac{Tn}{p-Tn-1})\sum_{s=1}^{T}\|w^{*}_{s}\|^{2}$
$=\frac{2}{T}\frac{n}{p}(1+\frac{\frac{T}{2}n}{p-Tn-1})((T^{2}-T)\cos\theta+T)\|w^{*}_{1}\|^{2}-\frac{Tn}{p}(1+\frac{1}{T}+\frac{Tn}{p-Tn-1})\|w^{*}_{1}\|^{2}$
Now, solving $\mathbb{E}[K(w_{1T})]=0$ yields
$\frac{2}{T}(1+\frac{\frac{T}{2}n}{p-Tn-1})((T-1)\cos\theta+1)=(1+\frac{1}{T}+\frac{Tn}{p-Tn-1})$
$\Rightarrow(2+\frac{Tn}{p-Tn-1})((T-1)\cos\theta+1)=(T+1+\frac{T^{2}n}{p-Tn-1})$
$\Rightarrow(2+\frac{Tn}{p-Tn-1})(T-1)\cos\theta=(T-1+\frac{T(T-1)n}{p-Tn-1})$
$\Rightarrow(2p-Tn-2)\cos\theta=p-1$
$\Rightarrow\cos\theta=\frac{p-1}{2p-Tn-2}$
C.3 Proof of Theorem B.1 and 4.1
To provide proof for both theorems, we start by presenting a more general expression and derive both theorems as special cases. We start by finding the loss of task $i$ at timestep $t$, i.e. $\mathbb{E}[\mathcal{L}_{i}(w_{{\bm{v}}{t}})]$, using an intermediate result in the proof provided in section C.2.2. In fact, we refer to Equation 22 and assume $n_{1}=n_{2}=...=n_{t-1}=m$, $n_{t}=n$, $n_{t+1}=...=n_{T}=0$ and denote $\bar{n}_{t}=(t-1)m+n$ to achieve:
@@
Proof of Theorem B.1
Proof.
Set $m=0$ in Equation 23, to retrieve the loss for the naive sequential learner:
$\mathbb{E}[\mathcal{L}_{i}(w_{t})]$ $=\frac{n}{p}\|w^{*}_{t}-w_{i}^{*}\|^{2}+(1-\frac{n}{p})\|w_{i}^{*}\|^{2}+\frac{n\sigma^{2}}{p-n-1}$
Therefore,
$\mathbb{E}[G(w_{T})]$ $=\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}[\mathcal{L}_{i}(w_{T})]$
$=\frac{n}{Tp}\sum_{i=1}^{T-1}\|w^{*}_{T}-w_{i}^{*}\|^{2}+\frac{1}{T}(1-\frac{n}{p})\sum_{i=1}^{T}\|w_{i}^{*}\|^{2}+\frac{n\sigma^{2}}{p-n-1}$
and for the forgetting, we have:
$\mathbb{E}[F(w_{T})]$ $=\frac{1}{T-1}\sum_{i=1}^{T}\mathbb{E}[\mathcal{L}_{i}(w_{T})]-\mathbb{E}[\mathcal{L}_{i}(w_{i})]$
$=\frac{n}{(T-1)p}\sum_{i=1}^{T-1}\|w^{*}_{T}-w_{i}^{*}\|^{2}+\frac{1}{T-1}(1-\frac{n}{p})\sum_{i=1}^{T}\|w_{i}^{*}\|^{2}+\frac{n\sigma^{2}}{p-n-1}$
$-\frac{1}{T-1}(1-\frac{n}{p})\sum_{i=1}^{T}\|w_{i}^{*}\|^{2}-\frac{n\sigma^{2}}{p-n-1}$
$=\frac{n}{(T-1)p}\sum_{i=1}^{T-1}\|w^{*}_{T}-w_{i}^{*}\|^{2}$
Proof of Theorem 4.1
Proof.
We set $T=2$ in Equation 23 to get:
@@
For notation simplicity, denote $r:=t+1$. Using Lemma C.6 and the notation introduced in Section C.1, we know that:
$w_{{\bm{v}}{t+1}}=(I-P_{1r})w_{{\bm{v}}{t}}+X^{\dagger}_{1r}y_{1r}$
where $P_{1r}=X_{1r}(X^{T}_{1r}X_{1r})^{-1}X^{T}_{1r}$ and $X^{\dagger}_{1r}=X_{1r}(X^{T}_{1r}X_{1r})^{-1}$. Notice that $X_{1r}\in\mathbb{R}^{p\times(tm+n)}$ is the result of the concatenation of all data points in the memory to the new data for task $t+1$. Importantly, $X_{1r}$ contains exactly $m$ columns from task $t$. Next we can write:
$w_{{\bm{v}}{t+1}}=(I-P_{1r})w_{{\bm{v}}{t}}+\sum_{s=1}^{r}P_{0s}w_{s}^{*}+X^{\dagger}_{1r}z_{1r}$
where $P_{0s}=X_{1r}(X^{T}_{1r}X_{1r})^{-1}X_{0s}$.Therefore,
@@
$=\mathbb{E}[\|\sum_{s=1}^{r}P_{0s}(w_{s}^{*}-w^{*}_{i})+(I-P_{1r})(w_{{\bm{v}}{t}}-w^{*}_{i})+X^{\dagger}_{1r}z_{1r}\|^{2}]$ $(\sum_{s=1}^{r}P_{0s}=P_{1r})$
$=\mathbb{E}[\|\sum_{s=1}^{r}P_{0s}(w_{s}^{*}-w^{*}_{i})+(I-P_{1r})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]+\mathbb{E}[\|X^{\dagger}_{1r}z_{1r}\|^{2}]$ $(z_{1r}\operatorname{\perp\!\!\!\perp}X_{1r},\mathbb{E}[z_{1r}]=0)$
$=\mathbb{E}[\|\sum_{s=1}^{r}P_{0s}(w_{s}^{*}-w^{*}_{i})\|^{2}]+\mathbb{E}[\|(I-P_{1r})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]+\mathbb{E}[\|X^{\dagger}_{1r}z_{1r}\|^{2}]$ $((I-P_{1r})P_{0s}=0)$
@@
The form of this expression is very similar to what we used in the proof of Theorem 3.1 in Section C.2.2. The only considerable difference that causes complications is the term $\mathbb{E}[\|(I-P_{1r})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$. The reason is that the term $w_{{\bm{v}}{t}}$ is no longer independent of $P_{1r}$, and it is not straightforward to calculate this expectation. Therefore, we avoid repeating the other steps and only focus on finding this term:
$\mathbb{E}[\|(I-P_{1r})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]=\mathbb{E}[\|w_{{\bm{v}}{t}}-w^{*}_{i}\|^{2}]-\mathbb{E}[\|P_{1r}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$
@@
Notice that $w_{{\bm{v}}{t}}$ is independent of $X_{r}$. Therefore, we decompose the second term using Lemma C.3 part (i):
$\mathbb{E}[\|P_{1r}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$ $=\mathbb{E}[\|P_{1t}(w_{{\bm{v}}{t}}-w^{*}_{i})+(I-P_{1t})X_{r}(X_{r}^{T}(I-P_{1t})X_{r})^{-1}X_{r}^{T}(I-P_{1t})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$
$=\mathbb{E}[\|P_{1t}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]+\mathbb{E}[\|(I-P_{1t})X_{r}(X_{r}^{T}(I-P_{1t})X_{r})^{-1}X_{r}^{T}(I-P_{1t})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$
@@
Recall that $X_{r}\in\mathbb{R}^{p\times n}$ follows the standard normal distribution and is independent of $(w_{{\bm{v}}{t}}-w^{*}_{i})$ and $P_{1t}$. Due to the rotational invariance property, we can write:
$\mathbb{E}[\|P_{1r}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$ $=\mathbb{E}[\|P_{1t}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]+(\frac{n}{p-tm})\mathbb{E}[\|(I-P_{1t})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$
$=(1-\frac{n}{p-tm})\mathbb{E}[\|P_{1t}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]+(\frac{n}{p-tm})\mathbb{E}[\|w_{{\bm{v}}{t}}-w^{*}_{i}\|^{2}]$
@@
Therefore, it suffices to focus on $\mathbb{E}[\|P_{1t}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$ as the next step. We go back to the definition of $w_{{\bm{v}}{t}}$ and use Lemma C.6 again. Notice that $w_{{\bm{v}}{t}}$ was trained on the samples in the memory from tasks $1$ to $t-1$ (each task with size $m$) in addition to the data from task $t$ with size $n$. Let’s build a new matrix $\hat{X}_{1t}\in\mathbb{R}^{p\times((t-1)m+n)}$ to represent all of the training data used in that step. Similarly, build $\hat{y}_{1t}$ and $\hat{z}_{1t}$ by concatenating all the label and noise values. Also denote $\hat{X}_{0s}\in\mathbb{R}^{p\times((t-1)m+n)}$ with zeros at all columns except the columns corresponding to task $s$. Then we can write:
$w_{{\bm{v}}{t}}=(I-\hat{P}_{1t})w_{{\bm{v}}{t-1}}+\sum_{s=1}^{t}\hat{P}_{0s}w_{s}^{*}+\hat{X}^{\dagger}_{1t}\hat{z}_{1t}$

### v2

@@ A.2 Proof of Theorems
underparameterized regime. In the underparameterized regime where $n_{t}\geq p+2$, the signle-task learner is the unique solution to the Equation 4 and is obtained by $w_{t}=(X_{t}X_{t}^{\top})^{-1}X_{t}y_{t}$ according to Lemma A.5. Therefore, we must have
$w_{t}=(X_{t}X_{t}^{\top})^{-1}X_{t}y_{t}=(X_{t}X_{t}^{\top})^{-1}X_{t}(X^{\top}_{t}w^{*}_{t}+z_{t})=w^{*}_{t}+(X_{t}X_{t}^{\top})^{-1}X_{t}z_{t}$
$\Rightarrow\mathbb{E}[\mathcal{L}_{t}(w)]=\mathbb{E}[\|w_{t}-w^{*}_{t}\|^{2}]=\mathbb{E}[\|(X_{t}X_{t}^{\top})^{-1}X_{t}z_{t}\|^{2}]$
@@
$\mathbb{E}[\mathcal{L}_{t}(w)]$ $=\mathbb{E}[z^{\top}_{t}X^{\top}_{t}(X_{t}X_{t}^{\top})^{-1}(X_{t}X_{t}^{\top})^{-1}X_{t}z_{t}]$
$=\sigma^{2}\mathbb{E}[\operatorname{tr}(X^{\top}_{t}(X_{t}X_{t}^{\top})^{-2}X_{t})]$
$=\sigma^{2}\mathbb{E}[\operatorname{tr}((X_{t}X_{t}^{\top})^{-2}X_{t}X^{\top}_{t})]$
$=\sigma^{2}\mathbb{E}[\operatorname{tr}((X_{t}X_{t}^{\top})^{-1})]$
$=\sigma^{2}\operatorname{tr}(\mathbb{E}[(X_{t}X_{t}^{\top})^{-1}]))$
@@
where the two last line comes from the fact that $(X_{t}X_{t}^{\top})^{-1}$ follows the inverse-Wishart distribution as $\mathcal{IW}(I_{p\times p},n)$.
overparameterized regime. In the overparameterized regime where $p\geq n_{t}+2$, we look for the solution of optimization 5. Therefore, based on Lemma A.6, we can have
@@
$=\mathbb{E}[\|(I-P_{t})w^{*}_{t}\|^{2}]+\mathbb{E}[z^{\top}_{t}(X^{\top}_{t}X_{t})^{-1}z_{t}]\qquad\quad(\mathbb{E}[z_{t}]=0)$
$=\mathbb{E}[\|(I-P_{t})w^{*}_{t}\|^{2}]+\frac{n_{t}\sigma^{2}}{p-n_{t}-1}\qquad\qquad\qquad((X^{\top}_{t}X_{t})^{-1}\sim\mathcal{IW}(I_{n_{t}\times n_{t}},p))$
@@
Based on Lemma A.6 we can write
$w_{1:T}=X_{1:T}(X^{\top}_{1:T}X_{1:T})^{-1}y_{1:T}$
@@
$\mathbb{E}[\mathcal{L}_{i}(w_{1:T})]=\mathbb{E}[\|w_{1:T}-w^{*}_{i}\|^{2}]$
$=\mathbb{E}[\|\sum_{s=1}^{T}P_{0|s}w^{*}_{s}+X_{1:T}^{\dagger}z_{1:T}-w^{*}_{i}\|^{2}]$
$=\mathbb{E}[\|\sum_{s=1}^{T}P_{0|s}(w^{*}_{s}-w_{i}^{*})+(I-P_{1:T})w_{i}^{*}+X_{1:T}^{\dagger}z_{1:T}\|^{2}]$
$=\mathbb{E}[\|\sum_{s=1}^{T}P_{0|s}(w^{*}_{s}-w_{i}^{*})+(I-P_{1:T})w_{i}^{*}\|^{2}]+\mathbb{E}[\|X_{1:T}^{\dagger}z_{1:T}\|^{2}]\qquad\quad(z_{1:T}\operatorname{\perp\!\!\!\perp}X_{1:T},\mathbb{E}[z_{1:T}]=0)$
$=\mathbb{E}[\|\sum_{s=1}^{T}P_{0|s}(w^{*}_{s}-w_{i}^{*})+(I-P_{1:T})w_{i}^{*}\|^{2}]+\sigma^{2}\operatorname{tr}(\mathbb{E}[(X_{1:T}X_{1:T})^{-1}])\qquad(z_{1:T}\sim\mathcal{N}(0,\sigma^{2}I))$
$=\mathbb{E}[\|\sum_{s=1}^{T}P_{0|s}(w^{*}_{s}-w_{i}^{*})+(I-P_{1:T})w_{i}^{*}\|^{2}]+\frac{\bar{n}\sigma^{2}}{p-\bar{n}-1}\quad\quad((X_{1:T}X_{1:T})^{-1}\sim\mathcal{IW}(I_{\bar{n}\times\bar{n}},p))$
@@
$\mathbb{E}[\|\sum_{s=1}^{T}P_{0|s}(w^{*}_{s}-w_{i}^{*})+(I-P_{1:T})w_{i}^{*}\|^{2}]$
$=\mathbb{E}[\|\sum_{s=1}^{T}P_{0|s}(w^{*}_{s}-w_{i}^{*})\|^{2}]+\mathbb{E}[\|(I-P_{1:T})w_{i}^{*}\|^{2}]$
@@
$=\sum_{s=1}^{T}\mathbb{E}[\|P_{0|s}(w^{*}_{s}-w_{i}^{*})\|^{2}]+\sum_{s=1}^{T}\sum_{\begin{subarray}{c}s^{\prime}=1\\ s^{\prime}\neq s\end{subarray}}^{T}\mathbb{E}[(w^{*}_{s}-w_{i}^{*})^{\top}P^{\top}_{0|s}P_{0s^{\prime}}(w^{*}_{s^{\prime}}-w_{i}^{*})]+(1-\frac{\bar{n}}{p})\|w_{i}^{*}\|^{2}$
Since the distribution of $P_{0|s}$ is invariant to the permutation of columns of $X_{1:T}$, we can focus on finding $\mathbb{E}[\|P_{0|1}(w^{*}_{1}-w_{i}^{*})\|^{2}]$ and $\mathbb{E}[(w^{*}_{1}-w_{i}^{*})^{\top}P^{\top}_{0|1}P_{0|T}(w^{*}_{T}-w_{i}^{*})]$ without loss of generality. In Lemmas A.7 and A.8, we have already calculated similar quantities. Therefore,
$\mathbb{E}[\|P_{0|1}(w^{*}_{1}-w_{i}^{*})\|^{2}]=\frac{n_{1}}{p}(1+\frac{\bar{n}-n_{1}}{p-\bar{n}-1})\|w^{*}_{1}-w_{i}^{*}\|^{2}$
@@
$\mathbb{E}[(w^{*}_{1}-w_{i}^{*})^{\top}P^{\top}_{0|1}P_{0|T}(w^{*}_{T}-w_{i}^{*})]=-\frac{n_{T}n_{1}}{p(p-\bar{n}-1)}\langle w^{*}_{1}-w_{i}^{*},w^{*}_{T}-w_{i}^{*}\rangle$
@@
$\mathbb{E}[G(w_{1:T})]$ $=\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}[\mathcal{L}_{i}(w_{1:T})]$
@@
$\mathbb{E}[K(w_{1:T})]$ $=\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}[\mathcal{L}_{i}(w_{i})]-\mathbb{E}[\mathcal{L}_{i}(w_{1:T})]$
@@
A.3 Proof of Theorem 4.1
Proof.
To provide the proof, we start by finding the loss of task $i$ at timestep $t$, i.e. $\mathbb{E}[\mathcal{L}_{i}(w_{{\bm{v}}{t}})]$, using an intermediate result in the proof provided in section A.2.2. In fact, we refer to Equation 21 and assume $n_{1}=n_{2}=...=n_{t-1}=m$, $n_{t}=n$, $n_{t+1}=...=n_{T}=0$ and denote $\bar{n}_{t}=(t-1)m+n$ to achieve:
@@
Next, we set $T=2$ in Equation 22 to get:
@@
For notation simplicity, denote $r:=t+1$. Using Lemma A.6 and the notation introduced in Section A.1, we know that:
$w_{{\bm{v}}{t+1}}=(I-P_{1:r})w_{{\bm{v}}{t}}+X^{\dagger}_{1:r}y_{1:r}$
where $P_{1:r}=X_{1:r}(X^{\top}_{1:r}X_{1:r})^{-1}X^{\top}_{1:r}$ and $X^{\dagger}_{1:r}=X_{1:r}(X^{\top}_{1:r}X_{1:r})^{-1}$. Notice that $X_{1:r}\in\mathbb{R}^{p\times(tm+n)}$ is the result of the concatenation of all data points in the memory to the new data for task $t+1$. Importantly, $X_{1:r}$ contains exactly $m$ columns from task $t$. Next we can write:
$w_{{\bm{v}}{t+1}}=(I-P_{1:r})w_{{\bm{v}}{t}}+\sum_{s=1}^{r}P_{0|s}w_{s}^{*}+X^{\dagger}_{1:r}z_{1:r}$
where $P_{0|s}=X_{1:r}(X^{\top}_{1:r}X_{1:r})^{-1}X_{0|s}$.Therefore,
@@
$=\mathbb{E}[\|\sum_{s=1}^{r}P_{0|s}(w_{s}^{*}-w^{*}_{i})+(I-P_{1:r})(w_{{\bm{v}}{t}}-w^{*}_{i})+X^{\dagger}_{1:r}z_{1:r}\|^{2}]$ $(\sum_{s=1}^{r}P_{0|s}=P_{1:r})$
$=\mathbb{E}[\|\sum_{s=1}^{r}P_{0|s}(w_{s}^{*}-w^{*}_{i})+(I-P_{1:r})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]+\mathbb{E}[\|X^{\dagger}_{1:r}z_{1:r}\|^{2}]$ $(z_{1:r}\operatorname{\perp\!\!\!\perp}X_{1:r},\mathbb{E}[z_{1:r}]=0)$
$=\mathbb{E}[\|\sum_{s=1}^{r}P_{0|s}(w_{s}^{*}-w^{*}_{i})\|^{2}]+\mathbb{E}[\|(I-P_{1:r})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]+\mathbb{E}[\|X^{\dagger}_{1:r}z_{1:r}\|^{2}]$ $((I-P_{1:r})P_{0|s}=0)$
@@
The form of this expression is very similar to what we used in the proof of Theorem 3.2 in Section A.2.2. The only considerable difference that causes complications is the term $\mathbb{E}[\|(I-P_{1:r})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$. The reason is that the term $w_{{\bm{v}}{t}}$ is no longer independent of $P_{1:r}$, and it is not straightforward to calculate this expectation. Therefore, we avoid repeating the other steps and only focus on finding this term:
$\mathbb{E}[\|(I-P_{1:r})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]=\mathbb{E}[\|w_{{\bm{v}}{t}}-w^{*}_{i}\|^{2}]-\mathbb{E}[\|P_{1:r}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$
@@
Notice that $w_{{\bm{v}}{t}}$ is independent of $X_{r}$. Therefore, we decompose the second term using Lemma A.3 part (i):
$\mathbb{E}[\|P_{1:r}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$ $=\mathbb{E}[\|P_{1:t}(w_{{\bm{v}}{t}}-w^{*}_{i})+(I-P_{1:t})X_{r}(X_{r}^{\top}(I-P_{1:t})X_{r})^{-1}X_{r}^{\top}(I-P_{1:t})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$
$=\mathbb{E}[\|P_{1:t}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]+\mathbb{E}[\|(I-P_{1:t})X_{r}(X_{r}^{\top}(I-P_{1:t})X_{r})^{-1}X_{r}^{\top}(I-P_{1:t})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$
@@
Recall that $X_{r}\in\mathbb{R}^{p\times n}$ follows the standard normal distribution and is independent of $(w_{{\bm{v}}{t}}-w^{*}_{i})$ and $P_{1:t}$. Due to the rotational invariance property, we can write:
$\mathbb{E}[\|P_{1:r}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$ $=\mathbb{E}[\|P_{1:t}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]+(\frac{n}{p-tm})\mathbb{E}[\|(I-P_{1:t})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$
$=(1-\frac{n}{p-tm})\mathbb{E}[\|P_{1:t}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]+(\frac{n}{p-tm})\mathbb{E}[\|w_{{\bm{v}}{t}}-w^{*}_{i}\|^{2}]$
@@
Therefore, it suffices to focus on $\mathbb{E}[\|P_{1:t}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$ as the next step. We go back to the definition of $w_{{\bm{v}}{t}}$ and use Lemma A.6 again. Notice that $w_{{\bm{v}}{t}}$ was trained on the samples in the memory from tasks $1$ to $t-1$ (each task with size $m$) in addition to the data from task $t$ with size $n$. Let’s build a new matrix $\hat{X}_{1:t}\in\mathbb{R}^{p\times((t-1)m+n)}$ to represent all of the training data used in that step. Similarly, build $\hat{y}_{1:t}$ and $\hat{z}_{1:t}$ by concatenating all the label and noise values. Also denote $\hat{X}_{0|s}\in\mathbb{R}^{p\times((t-1)m+n)}$ with zeros at all columns except the columns corresponding to task $s$. Then we can write:
$w_{{\bm{v}}{t}}=(I-\hat{P}_{1:t})w_{{\bm{v}}{t-1}}+\sum_{s=1}^{t}\hat{P}_{0|s}w_{s}^{*}+\hat{X}^{\dagger}_{1:t}\hat{z}_{1:t}$